1.) 2x^2 - 4x = y + 1
x + y = 1

Answer by ikleyn(52778)

.
1.) 2x^2 - 4x = y + 1
x + y = 1
~~~~~~~~~~~~~~~~~~~~`

Express y = 1-x from the second equation, and then substitute it into the first equation replacing y. You will get


2x^2 - 4x = (1-x) + 1,

2x^2 - 4x = 2 - x,

2x^2 - 3x - 2 = 0.

Continue from this point.


For many other similar solved problems see the lesson
    - Solving the system of algebraic equations of degree 2 and degree 1,
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of equations that are not linear".